OpenStudy (anonymous):

Solve the equation. Show your work. 1/a = 6/18

OpenStudy (anonymous):

K ! we have : 1/a=6/18 Right ? If we have : 2a = 16 What do u do to answer ?:) a=?

OpenStudy (anonymous):

@littleweezy

OpenStudy (anonymous):

isnt it 6/18 because if you get rid of the 1 then u end up with the 6/18!?!?!?!?!

OpenStudy (anonymous):

@AHazen : Step by step !:)